OpenStudy (anonymous):

WILL GIVE MEDAL FOR RIGHT ANSWER!!! Which ratios are proportional to 10/45 ? choose ALL the answers that are correct a 2/9 b 5/15 c 6/27 d 12/19

OpenStudy (anonymous):

more than one right answer

OpenStudy (catlover5925):

it is C....I took all of them...2 and divided it by 10 and 9 and divided it by 45...in order for it to be proportionate it needs to come out the same number when divided...the only one I got that came out with the same number was C

OpenStudy (anonymous):

c and?

OpenStudy (anonymous):

b and c

OpenStudy (anonymous):

It's only a

OpenStudy (anonymous):

10/2=5 45/9=5

OpenStudy (anonymous):

10/5=2 45/15=3 not proportioned

OpenStudy (catlover5925):

oh yes it A and C

OpenStudy (anonymous):

10/6= 1.666666667 45/27=1.666666667

OpenStudy (catlover5925):

10 divided by 6 = 1.666667 and 45 divided by 27 is 1.666667

OpenStudy (anonymous):

yes it's a and c

OpenStudy (anonymous):

yes @AhhhhhplzzzzzHelp is wright